OpenStudy (anonymous):

is this radical in its simplified form?: 9squarRt10/squarRt10?

OpenStudy (anonymous):

just need to know if i simplified it all i cud

OpenStudy (lgbasallote):

nope. not yet simplified

OpenStudy (lgbasallote):

you mean \[\huge \frac{9\sqrt{10}}{\sqrt{10}}\] right?

OpenStudy (anonymous):

yes